OpenStudy (anonymous):

Factor each trinomial below. Please show your work and check your answer. (1 point each) x2 – 8x + 15 a2 – a – 20 a2 – 5a – 20 a2 + 12ab + 27b2 2a2 + 30a + 100

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

For the first one: x^2 - 8 x + 15 x^2 - 5x - 3x + 15 x(x-5) -3(x-5) (x-5)(x-3) The check: x(x-3)-5(x-3) x^2-3x-5x+15 x^2-8x+15

OpenStudy (anonymous):

Yeah thats what I got I wasn't sure I actually got (x-3)(x-5) but its the same thing right?

OpenStudy (anonymous):

yep same thing

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

what did you get for the others and I'll tell you if they are correct

OpenStudy (anonymous):

or you haven't done them yet?

OpenStudy (anonymous):

for the second one I got (a+8)(a-5)

OpenStudy (anonymous):

no that's incorrect

OpenStudy (anonymous):

the second one should be (a-5)(a+4)

OpenStudy (anonymous):

Im working on them as we speak Im slow in math thats why its taking me a while to reply :/ math is my weakest subject

OpenStudy (anonymous):

Can you show me how I got that because Idk how I screwed up

OpenStudy (anonymous):

don't worry... speed isn't the key. as long as you do them properly and understand your steps

OpenStudy (anonymous):

the way to think about it is: first make sure there's a 1in front of your first exponent, for instance in the 2nd question: a2 – a – 20, the first exponent is a^2 which is also 1*a^2. There is a 1 in front of it

OpenStudy (anonymous):

And the 3rd one it came out to -4(a-5)

OpenStudy (anonymous):

ok I can understand that because the "a" is really just 1^2 right?

OpenStudy (anonymous):

the second step is multiply the first with the 3rd, i.e. a^2 and 20 = 20a^2 now, you're looking for two numbers which add to give you the second (-a), and multiply to give you 20a^2

OpenStudy (anonymous):

no, a is a. when there's no number in front of it, it's the same as 1 x a^2

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

yeah you want to begin by removing the number in front of your a^2. So if you had a 5a^2, you would divide all of the numbers by 5 so that you're left with a^2

OpenStudy (anonymous):

In this case, there's no number in front of a^2 so you're good

OpenStudy (anonymous):

anyhow, the two numbers that multiply to give you -20x, and add to give you -a are -5x and +4x

OpenStudy (anonymous):

so it would be (a-5)(a+4) because 5*4 = 20 and the 20 a's = a^2 so 20a^2?

OpenStudy (anonymous):

wouldn't it be (a+5)(a+4) because its a positive 20?

OpenStudy (anonymous):

because (-5x)*(+4x) = -20x (because - * + = -, and 5*4 = 20)

OpenStudy (anonymous):

oh ok that makes a lot more sense thanks!

OpenStudy (anonymous):

the question has a -20. sorry i said x instead of a :)

OpenStudy (anonymous):

yeah. so that's how you get (a-5)(a+4)

OpenStudy (anonymous):

so the work would come out to a^2-a-20 a^2*20 = 20a^2 (a-5)(a+4) ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you'd write it as a^2 -a -20 a^2 -5a+4a-20 a(a-5)+4(a-5) (a-5)(a+4)

OpenStudy (anonymous):

essentially, you break the center into the two values that you found

OpenStudy (anonymous):

ok thanks Im skipping the 3rd one and working on the 4th one right now because I just completely got lost in the 3rd one

OpenStudy (anonymous):

the 4th question is the same idea: a2 + 12ab + 27b2 you want to find the two values that will add to give you 12ab and multiply to give you 27a^2b^2

OpenStudy (anonymous):

27 being an odd number how would do this because 27 won't equal out to any 2 numbers

OpenStudy (anonymous):

sorry i had to log out for a second

OpenStudy (anonymous):

two values that add to give +12ab and multiply to give +27a^2b^2 is +9ab and +3ab

OpenStudy (anonymous):

a2 + 12ab + 27b2 a2 + 9ab + 3ab + 27b2 a(a+9b) + 3b(a+9b) (a+9b)(a+3b)

OpenStudy (anonymous):

are you still there?

OpenStudy (anonymous):

And the fifth one is: 2a2 + 30a + 100 2(a^2 + 15a + 50) the values that add to give 15a, and multiply to 50a^2 are 10a and 5a 2(a^2 + 5a + 10a + 50) 2[a(a+5) +10(a+5)] 2(a+5)(a+10)

OpenStudy (anonymous):

Thank you very much i was working trying to work on the second part of the paper right now thats why I didn't answer but thanks so much you have made my life a lot easier. the second part is a lot harder from the first thats why I skipped ahead and went onto it It’s your turn to be a game show host! As you know, in the game of Math Time, the contestants are given an answer and they must come up with the question that corresponds to the given answer. Your task for this portion of the assignment is to create two different “answers” (and the questions that accompany them) that the host could use for the final round of Math Time. The questions and answers you create must be unique. Check out the example and hint below, if needed. - second part the example it gives me is i.e. Question for Host: x2 + 6x + 5 is the product of these two binomials. Expected Question from Contestant: What is (x + 5)(x + 1)?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

looks fun enough... it's exactly the same thing as the first set. they give you a binomial and you have to factor it to get the question

OpenStudy (anonymous):

So would x^2 + 11x + 10 what is (x+10)(x+1) be a correct answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you!